# Decomposing God Header File via Multi-View Graph Clustering

### v1

@@ Abstract
God Header File refers to a header file with large code size and wide file impact. Such files pose difficulties in code comprehension and slow down compilation, ultimately increasing the maintenance cost during software evolution. Although this concept is similar to God Class, existing refactoring methods for God Classes are inappropriate for God Header Files. The reason lies in the fact that the code elements in header files are mostly short declaration types, and build dependencies of the entire system should be considered with the aim of improving compilation efficiency. Meanwhile, these methods overlook the concern of cyclic dependencies, which holds immense importance in the God Header File decomposition. To address these challenges, this paper proposes a God Header File decomposing approach based on multi-view graph clustering. It first constructs a code element graph with multiple relationships. Then after coarsening the graph, a novel multi-view graph clustering algorithm is applied to identify clusters of closely related code elements, and a heuristic algorithm is introduced to address the cyclic dependencies in the clustering result. We evaluate our approach on a synthetic dataset as well as six real-world God Header Files from different projects. The results show that our approach could achieve 11.5% higher accuracy in comparison to existing God Class refactoring methods. Moreover, our decomposition results attain better modularity on all the real-world God Header Files and reduce recompilation time for historical commits by 15% to 60%.
@@
software maintenance, code refactoring, header file, software re-engineering
@@ I Introduction
Code refactoring plays a crucial role in long-lifespan software projects. As the software is enhanced, modified, and adapted to new requirements, the code becomes more complex and drifts away from its original design, thereby lowering the quality of the software [1]. Refactoring enables developers to transform poorly designed or even chaotic code into well-structured one [2], thus improving the software’s robustness, reusability, performance, and other essential attributes.
One of the widely recognized code refactoring tasks is God Class refactoring. God Class [3] refers to a large class with many responsibilities in a system, which poses difficulties for developers in code comprehension, testing, and maintenance [4]. To deal with God Classes, some approaches have been proposed to refactor them by extracting new classes automatically, including static analysis-based methods [5, 6, 7], metric-based methods [8, 9] and deep learning-based methods [10]. These approaches pay attention to optimizing software structure, resulting in refactored classes with higher cohesion and lower coupling.
In our collaboration with an embedded software development enterprise, we encountered a similar situation, God Header File, where some header files exhibit large code size and wide file impact in C language projects. These header files tend to consolidate multiple responsibilities of the system, contain lots of declarations and are included by numerous code files. Once such header files are modified, all files that include them, which account for a large proportion of the entire software, must be recompiled, even if the modification is unrelated to these files. Fig 1 shows an example of such header files. The file guc.h is a header file from a cloud-native database software project PolarDB-for-PostgreSQL [11]. It contains 63 macro definitions, 19 data structure definitions, and 493 declarations of variables or functions. Any change to it triggers recompilation of 388 files directly including it (such as hooks.c) or transitively including it through other header files (such as tablecmds.c and dml.c), totaling 741,265 lines of code, incurring a significant recompilation cost. Since such header files pose challenges for comprehension and maintenance, it’s necessary to refactor them so as to optimize software structure and decrease compilation cost during software evolution.
Existing Methods and Limitations. Though we could draw inspiration from existing God Class refactoring methods[12], they are not suitable for decomposing God Header Files due to the following three reasons. First, existing approaches primarily focus on the internal dependencies among methods in a God Class, but we have to take into account build dependencies[13] that lie outside the God Header File in order to improve compilation efficiency. Second, existing God Class refactoring methods usually combine multiple code relationships through weighted summation. However, determining the weights for each type of relationship is quite challenging and the weights might vary across different projects [7]. Last, these methods have overlooked the concern of cyclic dependencies among decomposed files, which is a kind of architectural anti-patterns [14] and particularly intolerant in header files as they may lead to compilation errors.
@@
Our Approach. To address the above challenges, we propose a God Header File decomposing method based on multi-view clustering that leverages different types of code relationships. First, we construct a code element graph considering not only the internal dependency and semantic relationships among code elements in a God Header File, but also the co-usage relationships, which reflects build dependencies of the entire project. Then, we coarsen the graph based on dependency relationships and employ a multi-view graph clustering algorithm to cluster code elements based on their semantic and co-usage relationships. Finally, a heuristic algorithm is introduced to address cyclic dependencies in the clustering result.
@@
Evaluation and Results Summary. We evaluate our approach on both synthetic God Header Files and real-world God Header Files. The real-world God Header Files are derived from a preliminary study on 557 projects on GitHub. The experimental results reveal that: 1) our approach achieves 11.5% higher accuracy in comparison to existing methods, and exhibits more consistent performance across different projects; 2) when applied to real-world God Header Files, our approach provides decomposition results with better modularity and acyclic dependencies; 3) our approach can reduce the recompilation cost for historical commits by 15% to 60%.
Contributions. This paper makes aware of the problem of God Header Files and gives the following contributions:
  • A God Header File decomposing approach via multi-view graph clustering, which achieves better results than existing methods across different projects.
@@
  • A preliminary study on God Head Files in open-source community (GitHub) , which shows God Header Files are widespread in open-source C language projects.
  • Our evaluation setup and results are publicly available to facilitate reproducibility and future work in this area[15].
II Preliminary Study and Background
Since there is an absence of empirical evidence on God Header Files, we conduct a preliminary study to investigate the prevalence of God Header Files in open-source community. Then we introduce the background about multi-view graph clustering.
II-A God Header Files
@@
In this section, we collected numerous open-source C language projects and analyzed their header files. We also manually annotated some samples to estimate the spread of God Header Files across the entire dataset. The study includes the following steps:
@@
Projects Selection. We collected C language projects with size of more than 10,000 KB and stars of more than 500 from GitHub, as these criteria indicate projects of higher quality. We eventually obtained 557 projects (downloaded by January 2023).
@@
Figure 2: Joint distribution of header files’ code size and file impact. Manually annotated header files are depicted as scatters.
Header Files Sampling and Annotation. We selected 6 well-documented projects for manual annotation (shown in Table I). We sampled header files from each project using a stratified approach: for each range defined by $i*100<$ code size $<(i+1)*100$ AND $j*0.1*100\%<$ file impact $<(j+1)*0.1*100\%$, where $i>1$ and $101<j<$ 10, we randomly selected one header file if available. This ensured a balanced sampling of files across varying ranges of code size and file impact. And a total of 59 samples were collected in this process. After that, two graduate students independently classified each sample as either a God Header File or a Non-God Header File. In case of disagreement, final adjudication was undertaken by the authors.
TABLE I: Projects for manual annotation and their typical God Header Files
@@
The annotation results are also depicted as scatter points in Figure 2, showing that files in the upper-right section of the plot are more likely to be God Header Files. To mitigate false positives, we regard header files with more than 400 code elements AND file impact exceeding 40% of the entire project’s code lines as God Header Files in our subsequent experiment. The thresholds we picked (dotted lines in Fig 2) are relatively high to ensure certainty in the identification of God Header Files, and could show the lower boundary of the prevalence of God Header Files, although may overlook some real ones.
As a result, we identified 649 God Header Files, and they came from 203 software projects. Namely, at least 37.5% of the projects in our dataset suffer from God Header Files. Among them, 103 files had more than 100 commits and 14 files had been modified more than 5,000 times. Their large code size, wide file impact and frequent modifications increase the burden for software maintenance and evolution, underscoring the pressing need for a dedicated God Header File refactoring method. All the data used in this preliminary study is available in our replication package [15].
Findings: God Header Files are widespread in open-source C language projects. Totally 649 God Head Files appear in 203 open-source projects, which means about 37.5% of the projects are affected by God Header Files.
II-B Multi-View Graph Clustering
As mentioned above, existing God Class refactoring methods usually combine multiple code relationships through weighted summation. Since the optimal weights are not consistent [7], they show varying performance across different projects. We face similar challenges in our task, but the emergence of multi-view graph clustering algorithms[22, 23, 24, 25] proves highly suitable for modularization-related software refactoring. By leveraging the consistency and complementary of different views, these algorithms exhibit superior effectiveness and generalization compared to single-view clustering
In our work, we employ a novel multi-view graph clustering algorithm, DuaLGR[25], as our clustering method for three reasons. First, software graphs often encompass non-homophilous edges, as entities from different modules can have connections.This algorithm performs well in both homophilous and low homophilous graphs thanks to its refinement process. Second, it addresses the inconsistency in optimal weights across projects by dynamically assigning weights and orders in the fusion process. Last, its scalability allows easy incorporation of additional relationships,, such as co-evolution, if deemed necessary.
@@ III Approach
Fig 3 presents an overview of the proposed approach that aims to automatically decompose a God Header File into several sub-header files. It can be divided into four parts: graph construction, graph coarsening, multi-view graph clustering, and cyclic dependency fixing. In the graph construction phase (section III-A), we parse the God Header File and extract code elements along with their relationships to construct the code element graph. The code elements include macros, data types, and function declarations, which cannot be further decomposed. The code relationships include dependency, semantic, and co-usage. Then in the graph coarsening phase (section III-B), we use the scarce but influential dependency relationships to coarsen the code element graph, ensuring that closely related code elements will not be separated in subsequent processes. In the multi-view clustering part (section III-C), we fuse the information from semantic and co-usage relationships, which exhibits more intricate but relatively weaker connections, to partition the coarsened graph through a novel multi-view graph clustering algorithm. Finally, in the cyclic dependency fixing (section III-D), we address the cyclic dependencies in the clustering results using a heuristic search and then provide our solution.
@@ III-A Code Element Graph Construction
$\vspace{-5pt}d_{i\rightarrow j}=\begin{cases}\frac{1}{|Successor_{j}|}\ &{\rm if}\ v_{i}\ {\rm uses}\ v_{j},\\ 0\ &{\rm otherwise}.\end{cases}$ (1)
$A^{d}_{i,j}=max\{d_{i\rightarrow j},d_{j\rightarrow i}\}.\vspace{-5pt}$ (2)
@@
$A^{s}_{i,j}=\frac{|Word_{i}\cap Word_{j}|}{|Word_{i}\cup Word_{j}|}.$ (3)
@@
$\vspace{-5pt}A^{c}_{i,j}=\frac{|File_{i}\cap File_{j}|}{|File_{i}\cup File_{j}|}.$ (4)
@@ III-B Dependency-Based Graph Coarsening
As mentioned before, dependency relationships inherently convey stronger connections than the other two types of relationships. Intuitively, if two code elements exclusively depend on each other, they should be placed within the same sub-header file. However, when a code element is dependent upon many other code elements, it is less certain to reach the same conclusion. Thus, in this phase, we coarsen the code element graph by merging code elements exhibiting high dependency weights, ensuring the tightly connected code elements remain together throughout subsequent procedures.
@@ III-C Multi-View Graph Clustering
To carry out the clustering task, we apply a novel multi-view graph clustering algorithm DuaLGR [25] to the coarsened graph. DuaLGR takes a shared feature and adjacency matrices of different graph views as input, and refines them by extracting high-level view-common information. This process highlights the role of homophilous edges while reducing the influence of non-homophilous ones. Furthermore, it adaptively assigns weights and orders to various views to aggregate them into a consensus graph. Then, each node is embedded according to the consensus graph through a GNN-based encoder, facilitating subsequent clustering.
@@
$\overline{A}^{r}=\alpha(\frac{1}{od^{r}}\sum_{i=1}^{od^{r}}(A^{r})^{i})+\Omega,$ (5)
@@
To fuse graph views, DuaLGR utilizes the pseudo label to calculate a weight $w^{r}$ for each view. Finally, the global consensus graph $A$ can be obtained by $A=w^{s}\overline{A}^{s}+w^{c}\overline{A}^{c}$. The consensus graph is then encoded by a Graph Convolution Network [26] and used for clustering with the K-means algorithm [27].
After the clustering process, we obtain several clusters of the coarsened graph. If a node in the coarsened graph represents multiple code elements in the original code element graph, all of them are considered to be in the same cluster to which the node belongs.
@@ III-D Cyclic Dependency Fixing
In the previous stage, we obtained a set of disjoint clusters of code elements $C=\{C_{1},\cdots,C_{K}\}$, each representing a sub-header file. If any code element in cluster $C_{i}$ relies on a code element in cluster $C_{j}$ (forming a dependency relationship), an inclusion relationship is established between $C_{i}$ and $C_{j}$. If such inclusion relationships forms cyclic dependencies, the decomposition result is unacceptable in our task. Therefore, our focus during this stage is on the resolution of cyclic dependencies.
$\begin{split}gain=\sum_{\begin{subarray}{c}v_{r}\in V_{m},\\ v_{s}\in C_{j}\end{subarray}}(A^{d}_{r,s}+A^{d}_{s,r})-\sum_{\begin{subarray}{c}v_{r}\in V_{m},\\ v_{s}\in C_{i}-V_{m}\end{subarray}}(A^{d}_{r,s}+A^{d}_{s,r})-|V_{m}|.\end{split}$ (6)

### v2

@@ Abstract
God Header Files, just like God Classes, pose significant challenges for code comprehension and maintenance. Additionally, they increase the time required for code recompilation. However, existing refactoring methods for God Classes are inappropriate to deal with God Header Files because the code elements in header files are mostly short declaration types, and build dependencies of the entire system should be considered with the aim of improving compilation efficiency. Meanwhile, ensuring acyclic dependencies among the decomposed sub-header files is also crucial in the God Header File decomposition. This paper proposes a multi-view graph clustering based approach for decomposing God Header Files. It first constructs and coarsens the code element graph, then a novel multi-view graph clustering algorithm is applied to identify the clusters and a heuristic algorithm is introduced to address the cyclic dependencies in the clustering results. To evaluate our approach, we built both a synthetic dataset and a real-world God Header Files dataset. The results show that 1) Our approach could achieve 11.5% higher accuracy than existing God Class refactoring methods; 2) Our decomposition results attain better architecture on real-world God Header Files, evidenced by higher modularity and acyclic dependencies; 3) We can reduce 15% to 60% recompilation time for historical commits that require recompiling.
@@
software maintenance, code refactoring, header file
@@ I Introduction
Code refactoring plays a crucial role in long-lifespan software projects. As the software is enhanced, modified, and adapted to new requirements, the code becomes more complex and drifts away from its original design, thereby degrading the quality of the software [1]. Refactoring enables developers to transform poorly designed or chaotic code into well-structured one [2], thus improving the software’s robustness, reusability, performance, and other essential attributes.
During our collaboration with an embedded software development enterprise, we encounter a pressing code refactoring problem, named God Header File, where some header files exhibit large code size and wide file impact in the C language projects. Once such a header file is modified in a commit, all files that include it must be recompiled too. Eventually, the recompilation time of the related code commit becomes unbearable. Figure 1 shows an example of such header files. The file guc.h is a header file from a cloud-native database software project PolarDB-for-PostgreSQL [3]. It contains 63 macro definitions, 19 data structure definitions, and 493 declarations of variables or functions. Any change to it triggers recompilation of 388 files directly including it (such as hooks.c) or transitively including it through other header files (such as tablecmds.c and dml.c), amounting to a total of 741,265 lines of code, incurring significant recompilation cost. Therefore, it’s necessary to refactor these God Header Files so as to optimize software structure and reduce compilation cost during software evolution.
@@
To deal with God Header Files, we try to draw inspiration from the well-recognized tasks of God Class refactoring. God Class [4] refers to a large class with many responsibilities in a system, which poses difficulties for developers in code comprehension, test, and maintenance [5]. To deal with God Classes, some approaches have been proposed to refactor them by extracting new classes automatically, including static analysis-based methods [6, 7, 8], metric-based methods [9, 10] and deep learning-based methods [11]. These approaches pay attention to optimizing the software structure, resulting in refactored classes with higher cohesion and lower coupling.
However, we found that these God Class refactoring methods are not suitable for decomposing God Header Files due to the following three reasons. First, existing approaches primarily focus on the internal dependencies among methods in a God Class, but we have to take into account build dependencies [12] that lie outside the God Header File in order to improve compilation efficiency. Second, existing methods usually combine multiple code relationships through weighted summation. However, determining the weights for each type of relationship is quite challenging and the weights might vary across different projects [8]. Last, these methods have overlooked the concern of cyclic dependencies among decomposed files, which is a kind of architectural anti-pattern [13] and particularly intolerant in header files as they may lead to compilation errors.
@@
To address the above challenges, we propose a God Header File decomposing method based on multi-view clustering that leverages different types of code relationships. First, we construct a code element graph considering not only the internal dependency and semantic relationships among code elements in a God Header File, but also the co-usage relationships, which reflect build dependencies of the entire project. Then, we coarsen the graph based on dependency relationships and employ a multi-view graph clustering algorithm to cluster code elements based on their semantic and co-usage relationships. Finally, a heuristic algorithm is introduced to address the cyclic dependencies in the clustering result.
@@
We evaluate our approach on both synthetic God Header Files and real-world God Header Files. The real-world God Header Files are derived from a preliminary study of 557 projects on GitHub. The experimental results reveal that: 1) our approach achieves 11.5% higher accuracy in comparison to existing methods, and exhibits more consistent performance across different projects; 2) when applied to real-world God Header Files, our approach provides decomposition results with better modularity and acyclic dependencies; 3) our approach can reduce the recompilation cost for historical commits by 15% to 60%.
Compared with existing work, this paper gives the following contributions:
@@
  • Make aware of the problem of God Header Files. We carried out a study on the C language projects in GitHub, which shows that God Header Files are widespread in open-source projects.
  • Propose a God Header File decomposing approach via multi-view graph clustering, which achieves better accuracy and modularity than existing methods. Meanwhile, it can reduce 15% to 60% recompilation time for the related historical commits.
  • Build a benchmark to investigate the effectiveness of God Header File decomposition approaches, which is publicly available [14] to facilitate future work in this area.
II Preliminary Study on God Header Files
Here we conduct a preliminary study to investigate the prevalence of God Header Files in open-source projects.
II-A Study Setup
@@
To carry out the study, we collected the open-source C language projects and analyzed their header files. We manually annotated some samples to estimate the spread of God Header Files on GitHub. The study includes the following steps:
@@
Projects Selection. We collected C language projects with size of more than 10,000 KB and stars of more than 500 from GitHub. Then we obtained 557 projects (downloaded by January 2023).
@@
Figure 2: Joint distribution of header files’ code size and file impact.
TABLE I: Projects for manual annotation and their typical God Header Files.
@@
Header Files Sampling and Annotation. We selected 6 well-documented projects for manual annotation (shown in Table I). We sampled header files from each project using a stratified approach: for each range defined by $i*100<$ code size $<(i+1)*100$ AND $j*0.1*100\%<$ file impact $<(j+1)*0.1*100\%$, where $i>1$ and $101<j<$ 10, we randomly selected one header file if available. This ensured a balanced sampling of files across varying ranges of code size and file impact. And a total of 59 samples were collected in this process. After that, two graduate students independently classified each sample as either a God Header File or not. In cases of disagreement, final adjudication was undertaken by the authors. The annotation results are depicted as scatter points in Figure 2. The red circles represent God Header Files and the blue crosses are Non-God Header Files. It indicates that header files with larger code size wider file impact are more likely to be God Header Files.
God Header Files Mining. Based on the above findings, we regard header files with more than 400 code elements AND file impact exceeding 40% of the its project as God Header Files in our subsequent experiment. The thresholds we picked (dotted lines in Figure 2) are relatively high to ensure certainty in the identification of God Header Files, although may overlook some real ones.
II-B Study Results
Based on the above work, we identified 649 God Header Files from 203 projects. These files constitute less than 1% of the 761,999 total header files, which could be attributed to the strict thresholds we selected. As for projects, 203 out of the 541 software projects contain God Header Files, indicating that at least 37.5% of the projects in our study suffer from this issue. Specifically, 163 projects have 1 to 3 God Header Files. There are also 4 projects containing more than 20 God Header Files. Our further investigation revealed that these projects created duplicate header files to support multiple versions of hardware devices, contributing to the higher counts.
We also examined the commit histories of the 649 identified God Header Files. Among them, 103 files had more than 100 commits, and 14 files had been modified more than 5,000 times. Their large code size, wide file impact and frequent modifications increase the burden for software maintenance and evolution, underscoring the pressing need for a dedicated God Header File refactoring method.
In summary, we can conclude that God Header Files are widespread in open-source C language projects. Many of these files have undergone frequent modifications throughout their lifespan, incurring significant recompilation overhead. This highlights the need for researchers and developers to pay more attention to God Header File refactoring in the future. Based on the study, we also built a dataset of God Header Files, including all the data used in this preliminary study. It is available at [14].
@@ III Approach
Figure 3 presents an overview of the proposed approach that aims to automatically decompose a God Header File into several sub-header files. It can be divided into four parts: graph construction, graph coarsening, multi-view graph clustering, and cyclic dependency fixing. In the graph construction phase (section III-A), we parse the God Header File and extract code elements along with their relationships to construct the code element graph. The code elements are macros, data types, and function declarations, which cannot be further decomposed. The code relationships include dependency, semantic, and co-usage. Then in the graph coarsening phase (section III-B), we use the scarce but influential dependency relationships to coarsen the code element graph, ensuring that closely related code elements will not be separated in subsequent processes. In the multi-view clustering part (section III-C), we fuse the information from semantic and co-usage relationships, which exhibits more intricate but relatively weaker connections, to partition the coarsened graph through a novel multi-view graph clustering algorithm. Finally, in the cyclic dependency fixing (section III-D), we address the cyclic dependencies in the clustering results using a heuristic search algorithm and then provide our solution.
@@ III-A Code Element Graph Construction
$d_{i\rightarrow j}=\begin{cases}\frac{1}{|Successor_{j}|}\ &{\rm if}\ v_{i}\ {\rm uses}\ v_{j}\\ 0\ &{\rm otherwise}\end{cases}$ (1)
$A^{d}_{i,j}=max\{d_{i\rightarrow j},d_{j\rightarrow i}\}$ (2)
@@
$A^{s}_{i,j}=\frac{|Word_{i}\cap Word_{j}|}{|Word_{i}\cup Word_{j}|}$ (3)
@@
$A^{c}_{i,j}=\frac{|File_{i}\cap File_{j}|}{|File_{i}\cup File_{j}|}$ (4)
@@ III-B Dependency-Based Graph Coarsening
As mentioned before, dependency relationships are usually more sparse but inherently convey stronger connections than the other two types of relationships. Intuitively, if two code elements exclusively depend on each other, they should be placed within the same sub-header file. However, when a code element is dependent upon many other code elements, it is less certain to reach the same conclusion. Thus, in this phase, we coarsen the code element graph by merging code elements exhibiting high dependency weights, ensuring the tightly connected code elements remain together throughout subsequent procedures.
@@ III-C Multi-View Graph Clustering
To carry out the clustering task, we apply a novel multi-view graph clustering algorithm DuaLGR [21] to the coarsened graph. DuaLGR takes a shared feature and adjacency matrices of different graph views as input, and refines them by extracting high-level view-common information. Furthermore, it adaptively assigns weights and orders to various views to aggregate them into a consensus graph. Then, each node is embedded according to the consensus graph through a GNN-based encoder, facilitating subsequent clustering.
Compared to other clustering algorithms, DuaLGR proves highly suitable for decomposing related software refactoring tasks for several reasons: First, software graphs often encompass non-homophilous edges, as entities from different modules can have connections. This algorithm performs well in both homophilous and low homophilous graphs thanks to its refinement process. Second, it addresses the inconsistency in optimal weights across projects [8] by dynamically assigning weights and orders in the fusion process. Last, its scalability allows easy incorporation of additional relationships, such as co-evolution, if deemed necessary.
@@
$\overline{A}^{r}=\alpha(\frac{1}{od^{r}}\sum_{i=1}^{od^{r}}(A^{r})^{i})+\Omega$ (5)
@@
To fuse graph views, DuaLGR utilizes the pseudo label to calculate a weight $w^{r}$ for each view. Finally, the global consensus graph $A$ can be obtained by
$\vspace{-6pt}A=w^{s}\overline{A}^{s}+w^{c}\overline{A}^{c}$ (6)
The consensus graph is then encoded by a Graph Convolution Network [22] and used for clustering with the K-means algorithm [23].
After the clustering process, we acquire several clusters of the coarsened graph. If a node in the coarsened graph represents multiple code elements in the original code element graph, all of them are considered to be in the same cluster to which the node belongs.
@@ III-D Cyclic Dependency Fixing
In the previous stage, we obtain a set of disjoint clusters of code elements $C=\{C_{1},\cdots,C_{K}\}$, each representing a sub-header file. If any code element in cluster $C_{i}$ relies on a code element in cluster $C_{j}$ (forming a dependency relationship), an inclusion relationship is established between $C_{i}$ and $C_{j}$. If such inclusion relationships forms cyclic dependencies, the decomposition result is unacceptable in our task. Therefore, our focus during this stage is on the resolution of cyclic dependencies.
@@ 22            $\{C_{i_{1}},\cdots,C_{i_{l}}\}\leftarrow best\_clusters$;
$\begin{split}gain=\sum_{\begin{subarray}{c}v_{r}\in V_{m},\\ v_{s}\in C_{j}\end{subarray}}(A^{d}_{r,s}+A^{d}_{s,r})-\sum_{\begin{subarray}{c}v_{r}\in V_{m},\\ v_{s}\in C_{i}-V_{m}\end{subarray}}(A^{d}_{r,s}+A^{d}_{s,r})-|V_{m}|\end{split}$ (7)
The worst-case time complexity of our heuristic search algorithm is $O(2^{n})$. However, in our specific scenario, the number of sub-header files derived from a God Header File is relatively small in practice. Consequently, the iteration times remain manageable, ensuring that the cost of our approach is acceptable.